# Boter: Bootstrapping Knowledge Selection and Question Answering for Knowledge-based VQA

### v1

@@ 3.2. Selector and Answerer
Selector. After obtaining the Top-k knowledge documents using DPR for the $i$-th sample, we aim to choose $t$ most important knowledge documents from the retrieved documents. where $t$ is smaller than $k$. As shown in Fig. 2, we firstly use the frozen image encoder and Q-former to extract the image features $\mathbf{V}_{i}$, where these features are extracted once and then used by the Selector and the Answerer. Then image features $\mathbf{V}_{i}$ are fed into the independent fully-connected layer to obtain the visual embeddings $\mathbf{E}^{v}_{i}$. We concatenate the question, a retrieved knowledge document, and the Selection prompt ”Does the retrieved knowledge document provide the key information to help answer the question?” into one sentence $S$. Next, visual embeddings $\mathbf{E}^{v}_{i}$ and the text are concatenated and fed into the LLM (Flan-T5 (Chung et al., 2022) is adopted in our work). Last, we use the probability of generating the word ‘yes’ as the score of each retrieved knowledge document $P_{i,j}$, denoted as $s_{i,j}=\bm{LLM}(concat(\mathbf{E}^{v}_{i},S_{i}))$, and we select top-t documents $\hat{\mathcal{P}}_{i}=\{\hat{P}_{i,1},\hat{P}_{i,2},...,\hat{P}_{i,t}\}$ based on the scores. The Selector can be conceptualized as follows:
(2) $\hat{\mathcal{P}}_{i}==Selector(I_{i},Q_{i},\mathcal{P}_{i}),|\hat{\mathcal{P}}_{i}|=t$
@@
Answerer. After obtaining the selected knowledge documents, we aim to reason on the knowledge to answer questions. As shown in Fig. 2, we process the same image features to obtain the different visual embeddings $\mathbf{E}^{v}_{i}$ via the fully-connected layer of the Answerer. Next, we concatenate the question and the knowledge into one sentence $S^{\prime}$ using the template ”Question: {} Knowledge: {} Answer: ”. We concatenate the visual embeddings and the text, which are fed into the LLM with different LoRA parameters to get the answer. The model outputs corresponding answers based on different documents. The Answerer can be conceptualized as follows:
(3) $a_{i}=Answerer(I_{i},Q_{i},\hat{\mathcal{P}}_{i})$
@@ 3.3. Cycle Training
Answerer Training. We use Eq. 2 to get the selected knowledge documents $\hat{\mathcal{P}}_{i}$. The image $I_{i}$ is fed into the frozen ViT and Q-former to obtain the image features $\mathbf{V}_{i}$. We use the trainable $FC_{ans}$ layer to output the visual embeddings $\mathbf{E}^{v}_{ans,i}$. We concatenate the visual embedding, the question $Q_{i}$ and each selected knowledge document $\hat{P}_{i,j}$ to construct $t$ triplets for the sample, where $j={1,2,\dots,t}$. Then we finetune the Answerer with LoRA under the supervision of the ground truth answer $\mathcal{A}_{i}$:
(4) $\mathbf{E}^{v}_{ans,i}=FC_{ans}(\mathbf{V}_{i}),$
@@
Selector Training. We first use Eq. 3 to predict answers based on each retrieved knowledge document $P_{i,j}$. Then we assign pseudo labels to the retrieved documents according to model predictions and weak supervision labels (Luo et al., 2021; Lin and Byrne, 2022; Lin et al., 2023). We use ”yes” and ”no” as pseudo labels, where label a document as positive knowledge if Answerer can output the correct answer using that document and the document contains any of the answers in $\mathcal{A}_{i}$.
(5) $y_{i,j}=\begin{cases}\text{yes},\quad\textit{if}\ \ a_{i}=a_{i}^{*}\wedge P_{i,j}\ \ \text{contains an answer in}\ \ \mathcal{A}_{i}\\ \text{no},\quad else\end{cases}$
Table 1. Performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset. Knowledge Sources: ConceptNet (C); Wikipedia (W); Google Search (GS); Google Images (GI). The best result in the table is bolded. Compared with the previous SOTA method Prophet (Shao et al., 2023) using GPT-3, our method achieves 62.83% performance, with a performance improvement of 1.73%. The results demonstrate the effectiveness of our method.
Models Large Models $K_{train}$ $K_{test}$ Knowledge Resource Accuracy (%)
BAN+AN (Marino et al., 2019) - - - W 25.6
ConceptBERT (Gardères et al., 2020) - - - C 33.7
KRISP (Marino et al., 2021) - - - C+W 38.4
Visual Retriever-Reader (Luo et al., 2021) - 100 100 GS 39.2
MAVEx (Wu et al., 2022) - - - W+C + GI 39.4
TRiG(Ensemble) (Gao et al., 2022) T5-large (770M) 100 100 W 50.5
KAT(Single) (Gui et al., 2021) T5-large (770M) 40 40 W + GPT-3 (175B) 53.1
KAT(Ensemble) (Gui et al., 2021) T5-large (770M) 40 40 W + GPT-3 (175B) 54.4
RA-VQA (Lin and Byrne, 2022) T5-large (770M) 5 50 GS 54.5
REVIVE(Single) (Lin et al., 2022) T5-large (770M) 40 40 W+GPT-3 (175B) 56.6
REVIVE(Ensemble) (Lin et al., 2022) T5-large (770M) 40 40 W+GPT-3 (175B) 58.0
BLIP2 T5-XXL (Dai et al., 2023) BLIP2 T5-XXL (11B) - - BLIP2 54.7
PICa (Yang et al., 2022) GPT-3 (175B) - - GPT-3 (175B) 48.0
Flamingo (Alayrac et al., 2022) Flamingo (80B) - - Flamingo (80B) 57.8
PromptCap (Hu et al., 2022a) GPT-3 (175B) - - GPT-3 (175B) 60.4
Prophet (Shao et al., 2023) GPT-3 (175B) - - GPT-3 (175B)+MCAN 61.1
Boter(Ours) BLIP2 T5-XL (3B) 5 5 GS 62.83
@@
4. EXPERIMENTS
4.1. Experimental Setup
Dataset. We conducted extensive experiments on OK-VQA (Marino et al., 2019) to evaluate the effectiveness of our method. OK-VQA is a challenging open-domain knowledge-based VQA dataset that requires models to leverage various external knowledge sources to answer questions. The dataset contains 14,055 questions and 14,031 images, whereas the training set and testing set have 9k and 5k image-question pairs, respectively. Due to no knowledge base being provided for OK-VQA, we need to choose the proper knowledge base for the dataset. In this paper, we adopt Google Search Corpus (Luo et al., 2021) as the knowledge base which is collected in the websites using the Google Search API. The dataset has been proven to be effective by previous methods (Luo et al., 2021; Lin and Byrne, 2022; Lin et al., 2023). We do not evaluate our method on A-OKVQA (Schwenk et al., 2022) like other retrieved-based methods (Gui et al., 2021; Lin et al., 2022; Lin and Byrne, 2022; Lin et al., 2023), as it emphasizes visually-grounded reasoning rather than knowledge retrieval and selection.
Evaluation Metric. We use the standard VQA metric (Antol et al., 2015) to evaluate the performance of the model. Given the prediction of the question $a$ and the groudtruth answer set $\mathcal{A}$, the VQA accuracy is calculated as:
(7) $Accuracy(a,\mathcal{A})=\min(\frac{\#A(a)}{3},1),$
@@ 4.1. Experimental Setup
Implementation Details. In our experiment, we adopt BLIP2 T5-XL (3B) (Li et al., 2023) to initialize the Selector and Retriever. We freeze the image encoder and Q-former, with both the Selector and Retriever sharing the same visual module. We finetune the fully connected layer and use LoRA (Hu et al., 2022b) to train the LLM. We use the default huggingface-PEFT setting: r=8, lora_alpha=32, lora_dropout=0.1. We use Adam as the optimizer and set the batch size to 8. We use the warm-up strategy which trains the model with an initial learning rate of 1e-4 and warm-up factor of 0.05 for 1000 steps and then utilizes a cosine annealing learning strategy with an initial learning rate of 1e-4 and a final learning rate of 0 after 10 epochs. We use top-30 knowledge documents retrieved by a pre-trained DPR (Karpukhin et al., 2020) as candidates for Selector and use the selected top-5 documents from the 30 documents for the Answerer to train and infer, denoted as $K_{candidate}=30,K_{train}=5,K_{test}=5$. We use 2 Nvidia A800 GPUs (80G) for all experiments. And our codes will be released upon paper acceptance.
4.2. Overall Results
We demonstrate the performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset in the Tab. 1. From the results, we have the following observations: (1) Early models (BAN+AN (Marino et al., 2019), ConceptBERT (Gardères et al., 2020), KRISP (Marino et al., 2021), Visual Retriever-Reader (Luo et al., 2021), and MAVEx (Wu et al., 2022)) adopt the small visual-language model as backbones and inject various knowledge bases (ConceptNet, Wikipedia, Google Web Search and Google Images) into the model, achieving a VQA accuracy from 25.6% to 39.4%. Although these models use external knowledge to help produce answers, they have relatively low performance because of the weak knowledge reasoning ability of traditional backbones. (2) Recent methods (TRiG(Gao et al., 2022), KAT (Gui et al., 2021), RA-VQA (Lin and Byrne, 2022), and REVIVE (Lin et al., 2022)) introduce the large language model to improve the capacity for common-sense reasoning. They exploit T5-large (Chung et al., 2022) to reason on the knowledge for question answering. Compared with early methods, they have a higher accuracy ranging from 50.5% to 58.0%. (3) Some recent methods take a very large language model GPT-3 (Brown et al., 2020) as the implicit knowledge base for OK-VQA. PICa (Yang et al., 2022) is the first method to use GPT-3. It uses the style of in-context learning to prompt GPT-3 to answer questions directly. Inspired by PICa, some methods (Hu et al., 2022a; Shao et al., 2023) improve the performance of GPT-3 by generating question-guided captions and providing answer candidates, achieving higher VQA score on OK-VQA, about 61%. Meanwhile, some retrieved-based methods (Gui et al., 2021; Lin et al., 2022) use GPT-3 to generate implicit knowledge and combine them with the retrieved explicit knowledge, which improves the performance. (4) Recent large multimodal large language models have achieved high performance on OK-VQA without using external knowledge (e.g., Flamingo (Alayrac et al., 2022) (80B) has a zero-shot VQA score of 57.8%, BLIP2 T5-XXL (Dai et al., 2023) has a finetuning VQA score of 54.7%). (5) Our method aims to select key knowledge and answer questions utilizing the multimodal perception and reasoning ability of the multimodal large language model. By cycle training, we bootstrap Selector and Answerer to perform knowledge selection and question answering consistently. With a small number of selected knowledge documents, our method outperforms other state-of-the-art models, achieving an accuracy of 62.83%. It is worth noting that we do not use GPT-3 and we only train the 0.16% parameters of the model.
4.3. Ablation Study
@@
Effect of Selector. We conduct the ablation study to evaluate the effectiveness of Selector in our method. We show the results in Tab. 2. From the results, we can observe: (1) Our framework, leveraging key knowledge documents selected by the Selector, consistently outperforms the Answerer when using the same number of documents retrieved by DPR. We improve the performance by 2.14% and 1.88% with 1 and 5 test knowledge documents, compared to DPR-based retrieval. When using the randomly selected documents, the model performs worst. These results demonstrate that top-ranked knowledge documents based on DPR scores are not optimal for question answering and our key knowledge selection module can identify relevant documents for accurate question answering, ensuring the coherence of knowledge retrieval and question-answering processes. (2) With an increase in the number of test knowledge documents to 10, the performance of the model DPR-based retrieval almost has no improvement and our method’s performance decreases. This suggests that additional documents may introduce more irrelevant knowledge, adversely affecting question-answering accuracy.
Table 2. Ablation study on the Selector. We select 5 knowledge documents from top-30 knowledge candidates retrieved by DPR. ’DPR Score’ refers to selecting top-5 knowledge based on similarity scores. ’Random Selection’ means randomly selecting 5 knowledge documents from 30 candidate knowledge documents. ’Selector’ denotes choosing 5 key knowledge documents by the Selector.
$K_{train}$ $K_{test}$ Knowledge Selection Accuracy (%)
@@
5 10 Random Selection 52.11
5 10 DPR Score 60.93
5 10 Selector 62.17
Effect of Answerer. In Tab. 3, we present a comparison of our Answerer using different knowledge reasoning methods. The results show that the performance using the strategy of voting surpasses that of concatenating under different knowledge selection settings. We argue that directly combining all the knowledge documents into a lengthened document makes it difficult for the Answerer to reason on them, which is easily influenced by noisy information. In contrast, it is easier for the Answerer to reason on each document to predict the answer. Simple voting can choose the best answer.
Table 3. Ablation study on different knowledge reasoning methods of Answerer. ’Concatenating’ denotes that we combine the key knowledge documents into one sentence and feed it into the Answerer to predict the final answer. ’Voting’ means that we feed different key knowledge documents into the Answerer to predict different answers and choose the best answer based on majority voting.
Method Backbone Accuracy (%)
Concatenating BLIP2 (fine-tuned) w knowledge from DPR 59.11
@@
Concatenating Boter (Ours) w knowledge from Selector 62.06
@@
Effect of cycle training. To evaluate the effectiveness of our cycle-training method, we compare the method with the strategy of independent training of two modules. We finetune the Answerer with the knowledge documents retrieved by DPR as the baseline. Independent training means that we train the Selector and the Answerer respectively. Initially, we train the Answerer module utilizing knowledge documents retrieved by DPR. Subsequently, employing the trained Answerer, we generate answers for each retrieved knowledge document, thereby generating pseudo-labels for the retrieved knowledge. We then proceed to train the Selector module supervised by these pseudo-labels. Finally, we conduct finetuning of the Answerer once more, incorporating new knowledge documents selected by the Selector. The results in Tab. 4 show that the model with cycle training outperforms the model with iterative refinement by 3.81%. The VQA score of using independent training is even lower than the baseline. These results demonstrate that our cycle training method can effectively boost the Selector and Answerer each other, which makes the model find key knowledge documents and leverage the knowledge to answer questions.
Table 4. Ablation study on different training methods of our framework.
Methods Accuracy (%)
@@
Effect of different methods of pseudo-labeling. In Tab. 5, we compare the model performance with different methods of pseudo-labeling. When using the model predictions as guidance, the model has a VQA score of 62.31%. When adding the weak supervision as the guidance, the model’s VQA score increases from 62.31% to 62.83%. The results demonstrate that using weak supervision labels preserves potentially useful documents, aiding the Answerer in accurately answering questions.
Table 5. Ablation study on different methods of pseudo-labeling.
Model predictions Weak supervision labels Accuracy (%)
@@
Effect of key knowledge documents ranges and quantities. In Tab. 6, we evaluate key knowledge document selection using various numbers of candidate documents and selected documents. From the results, we have the following findings: (1) As the number of selected documents increases, the model’s performance improves. This indicates that using more documents to train and test contributes to answering questions. (2) Using more documents for training can improve the performance a lot (the 2nd line $v.s.$ the last line). However, using more documents for testing has almost no improvement (the 3rd line $v.s.$ 4th line). (3) When the number of candidate documents increases, the model’s performance improves. The result demonstrates that low-ranked documents based on DPR scores may contain useful information for question answering. It is necessary for the model to select key knowledge documents.
Table 6. Ablation study on different numbers of candidate documents and selected documents.
$K_{candidate}$ $K_{train}$ $K_{test}$ Accuracy (%)
@@
Effect of different vision information in our framework. In Tab. 7, we explore the impact of different vision information under different settings. We directly finetune BLIP2 without using knowledge and we find that the model using LLM-aligned query features outperforms the model using visual contexts, improving the accuracy from 41.82% to 55.11%. Using both can bring the performance improvement a little. In our framework, using query features improves the VQA score from 52.93% to 62.83%. These results demonstrate that introducing LLM-aligned visual features can help the model understand the visual details better, which contributes to improving the model’s performance. Using both degrades the performance a little.
Table 7. Ablation study on different vision information in our framework.
Models Query features Context Accuracy (%)
BLIP2 (fine-tuned) w/o knowledge $\checkmark$ 55.11
$\checkmark$ 41.82
$\checkmark$ $\checkmark$ 55.43
Boter (Ours) $\checkmark$ 62.83
$\checkmark$ 52.93
$\checkmark$ $\checkmark$ 62.52
Table 8. Ablation study on different documents selection in Answerer fine-tuning.
Knowledge Selection Accuracy (%)
@@
Figure 3. Qualitative results on the test split of OK-VQA. We compared our method with a model that fine-tunes BLIP2 with knowledge ranked by DPR. The middle segment of the graph represents knowledge from various methods used to answer questions. On the right side of the graph, different answers are depicted when using distinct knowledge. Green and red colors indicate whether the selected final answer is correct.
Effect of different knowledge documents selection in Answerer fine-tuning. Tab. 8 compares the Answerer fine-tuning with different document selection strategies. The results show that our framework performs optimally when utilizing Selector in both Answerer training and inference. This is likely because the Selector provides more informative key knowledge documents and using both Selector ensures the consistency between the training domain and testing domain.
4.4. Qualitative Analysis

### v3

@@ 3.2 Selector and Answerer
After obtaining the Top-k knowledge documents using DPR for the $i$-th sample, we aim to choose $t$ most important knowledge documents from the retrieved documents, where $t$ is smaller than $k$. As shown in Fig. 1, we first use the frozen image encoder and Q-former to extract the image features $\mathbf{V}_{i}$, where these features are extracted once and then used by the Selector and the Answerer. Then image features $\mathbf{V}_{i}$ are fed into the independent fully-connected layer to obtain the visual embeddings $\mathbf{E}^{v}_{i}$. We concatenate the question, a retrieved knowledge document, and the Selection prompt "Does the retrieved knowledge document provide the key information to help answer the question?" into one sentence $S$. Next, visual embeddings $\mathbf{E}^{v}_{i}$ and the text are concatenated and fed into the LLM (Flan-T5 Chung et al. (2022) is adopted in our work). Last, we use the probability of generating the word ‘yes’ as the score of each retrieved knowledge document $P_{i,j}$, denoted as $s_{i,j}=\bm{LLM}(concat(\mathbf{E}^{v}_{i},S_{i}))$, and we select top-t documents $\hat{\mathcal{P}}_{i}=\{\hat{P}_{i,1},\hat{P}_{i,2},...,\hat{P}_{i,t}\}$ based on the scores. The Selector can be conceptualized as follows:
$\hat{\mathcal{P}}_{i}==Selector(I_{i},Q_{i},\mathcal{P}_{i}),|\hat{\mathcal{P}}_{i}|=t$ (1)
Answerer.
@@
After obtaining the selected knowledge documents, we aim to reason on the knowledge to answer questions. As shown in Fig. 1, we process the same image features to obtain the different visual embeddings $\mathbf{E}^{v}_{i}$ via the fully-connected layer of the Answerer. Next, we concatenate the question and the knowledge into one sentence $S^{\prime}$ using the template "Question: {} Knowledge: {} Answer: ". We concatenate the visual embeddings and the text, which are fed into the LLM with different LoRA parameters to get the answer. The model outputs corresponding answers based on different documents. The Answerer can be conceptualized as follows:
$a_{i}=Answerer(I_{i},Q_{i},\hat{\mathcal{P}}_{i})$ (2)
@@ 3.3 Self-Bootstrap Learning
Answerer Training.
We use Eq. 1 to get the selected knowledge documents $\hat{\mathcal{P}}_{i}$. The image $I_{i}$ is fed into the frozen ViT and Q-former to obtain the image features $\mathbf{V}_{i}$. We use the trainable $FC_{ans}$ layer to output the visual embeddings $\mathbf{E}^{v}_{ans,i}$. We concatenate the visual embedding, the question $Q_{i}$ and each selected knowledge document $\hat{P}_{i,j}$ to construct $t$ triplets for the sample, where $j={1,2,\dots,t}$. Then we finetune the Answerer with LoRA under the supervision of the ground truth answer $\mathcal{A}_{i}$:
$\mathbf{E}^{v}_{ans,i}=FC_{ans}(\mathbf{V}_{i}),$ (3)
@@
Selector Training.
We first use Eq. 2 to predict answers based on each retrieved knowledge document $P_{i,j}$. Then we assign pseudo labels to the retrieved documents according to model predictions and weak supervision labels Luo et al. (2021); Lin and Byrne (2022); Lin et al. (2023). We use "yes" and "no" as pseudo labels, where label a document as positive knowledge if Answerer can output the correct answer using that document and the document contains any of the answers in $\mathcal{A}_{i}$.
$y_{i,j}=\begin{cases}\text{yes},\quad\textit{if}\ \ a_{i}=a_{i}^{*}\wedge\\ \quad\quad P_{i,j}\ \ \text{contains an answer in}\ \ \mathcal{A}_{i}\\ \text{no},\quad else\end{cases}$ (4)
Table 1: Performance comparison with state-of-the-art (SOTA) methods on the OK-VQA dataset. Knowledge Sources: ConceptNet (C); Wikipedia (W); Google Search (GS); Google Images (GI). The best result in the table is bolded. The results show that our method achieves the state-of-the-art performance.
Models Large Models $K_{train}$ $K_{test}$ Knowledge Resource Acc (%)
BAN+AN Marino et al. (2019) - - - W 25.6
ConceptBERT Gardères et al. (2020) - - - C 33.7
KRISP Marino et al. (2021) - - - C+W 38.4
Visual Retriever-Reader Luo et al. (2021) - 100 100 GS 39.2
MAVEx Wu et al. (2022) - - - W+C + GI 39.4
PICa Yang et al. (2022) GPT-3 (175B) - - GPT-3 48.0
TRiG(Ensemble) Gao et al. (2022) T5-large (770M) 100 100 W 50.5
KAT(Single) Gui et al. (2021) T5-large (770M) 40 40 W + GPT-3 53.1
KAT(Ensemble) Gui et al. (2021) T5-large (770M) 40 40 W + GPT-3 54.4
RA-VQA Lin and Byrne (2022) T5-large (770M) 5 50 GS 54.5
REVIVE(Single) Lin et al. (2022) T5-large (770M) 40 40 W+GPT-3 56.6
REVIVE(Ensemble) Lin et al. (2022) T5-large (770M) 40 40 W+GPT-3 58.0
PromptCap Hu et al. (2022b) GPT-3 (175B) - - GPT-3 60.4
Prophet Shao et al. (2023) GPT-3 (175B) - - GPT-3+MCAN 61.1
FillingGap Wang et al. (2023) GPT-3 (175B) - - GPT-3 61.3
SimpleBaseline Xenos et al. (2023) LLaMA 2 (13B) - - LLaMA 2 61.2
Cola-FT Chen et al. (2024) FLAN-T5(11B) - - BLIP+OFA 62.4
Flamingo Alayrac et al. (2022) Flamingo (80B) - - Pretrain 57.8
InstructBLIP Dai et al. (2023) InstructBLIP Vicuna (7B) - - Pretrain 62.1
Qwen-VL Bai et al. (2023) Qwen-VL(Qwen-7B) - - Pretrain 58.6
MM-Reasoner Khademi et al. (2023) Flamingo (80B) - - GPT-4 60.8
BLIP2 (fine-tuned) Li et al. (2023) BLIP2 T5-XL (3B) - - Pretrain 55.4
RA-VQA-v2 Lin et al. (2023) BLIP2 T5-XL (3B) 5 5 GS 62.1
PreFLMR Lin et al. (2024) BLIP2 T5-XL (3B) 5 5 GS 61.8
Ours BLIP2 T5-XL (3B) 5 5 GS 62.8
@@
4 Experiments
4.1 Experimental Setup
Dataset.
We conduct extensive experiments on OK-VQA Marino et al. (2019) to evaluate the effectiveness of our method. OK-VQA is a challenging open-domain knowledge-based VQA dataset that requires models to leverage various external knowledge sources to answer questions. The dataset contains 14,055 questions and 14,031 images, whereas the training set and testing set have 9k and 5k image-question pairs, respectively. Due to no knowledge base being provided for OK-VQA, we need to choose the proper knowledge base for the dataset. In this paper, we adopt Google Search Corpus Luo et al. (2021) as the knowledge base which is collected in the websites using the Google Search API.
Evaluation Metric.
We use the standard VQA metric Antol et al. (2015) to evaluate the performance of the model. Given the prediction of the question $a$ and the groudtruth answer set $\mathcal{A}$, the VQA accuracy is calculated as:
$Accuracy(a,\mathcal{A})=\min(\frac{\#A(a)}{3},1),$ (6)
@@ 4.1 Experimental Setup
In our experiment, we adopt BLIP2 T5-XL (3B) Li et al. (2023) to initialize the Selector and Answerer. We freeze the image encoder and Q-former, with both the Selector and Answerer sharing the same visual module. We finetune the fully connected layer and use LoRA Hu et al. (2022a) to train the LLM. We use the default huggingface-PEFT setting: r=8, lora_alpha=32, lora_dropout=0.1. We use Adam as the optimizer and set the batch size to 8. We use the warm-up strategy which trains the model with an initial learning rate of 1e-4 and warm-up factor of 0.05 for 1000 steps and then utilizes a cosine annealing learning strategy with an initial learning rate of 1e-4 and a final learning rate of 0 after 10 epochs. We use top-30 knowledge documents retrieved by a pre-trained DPR Lin and Byrne (2022) as candidates for Selector and use the selected top-5 documents from the 30 documents for the Answerer to train and infer, denoted as $K_{candidate}=30,K_{train}=5,K_{test}=5$. We use 2 Nvidia A800 GPUs (80G) for all experiments.
4.2 Comparison with State-of-the-art Methods
As shown in Tab. 1, we can see that early models (BAN+AN Marino et al. (2019), ConceptBERT Gardères et al. (2020), KRISP Marino et al. (2021), Visual Retriever-Reader Luo et al. (2021), and MAVEx Wu et al. (2022)) have a weak performance, achieving a VQA accuracy from 25.6% to 39.4%. Recently, by introducing larger models (T5-large, GPT-3, LLaMA, Vicuna) and diverse knowledge resources (ConceptNet, Wikipedia, Google Web Search and Google Images), the performance has a significant performance improvement, achieving a VQA accuracy of 62.4%. Our method aims to augment the reasoning ability to answer knowledge-intensive questions of the large visual-language model. When directly finetuning BLIP2 T5-XL on OKVQA, the model has a low performance of 55.44%. By introducing external knowledge, the performance has a significant performance improvement. Different from RA-VQA-v2 Lin et al. (2023) and PreFLMR Lin et al. (2024), we do not train a multimodal retriever from scratch which requires expensive annotations and high computational costs. We directly leverage the large visual-language model to select key knowledge from the retrieved knowledge by DPR like the process of re-ranking. With the same knowledge resources (i.e., Google Search), our method achieves 62.8% accuracy, outperforming other state-of-the-art models. It is worth noting that we do not use GPT-3 and we only train the 0.16% parameters of the model. These results demonstrate the effectiveness of the proposed approach.
4.3 Ablation Study
@@
Effect of Selector.
We conduct the ablation study to evaluate the effectiveness of Selector in our method. We show the results in Tab. 2. From the results, we can observe: our framework, leveraging key knowledge documents selected by the Selector, consistently outperforms the Answerer when using the same number of documents retrieved by DPR. We improve the performance by 2.14% and 1.88% with 1 and 5 test knowledge documents, compared to DPR-based retrieval. When using the randomly selected documents, the model performs worst. These results demonstrate that top-ranked knowledge documents based on DPR scores are not optimal for question answering and our key knowledge selection module can identify relevant documents for accurate question answering, ensuring the coherence of knowledge retrieval and question-answering processes.
Table 2: Comparison of our selector with different knowledge selection strategies. We select 5 knowledge documents from top-30 knowledge candidates retrieved by DPR. DPR Score refers to selecting top-5 knowledge based on similarity scores. Random Selection means randomly selecting 5 knowledge documents from 30 candidate knowledge documents. Selector denotes choosing 5 key knowledge documents by the Selector.
$K_{train}$ $K_{test}$ Knowledge Selection Acc (%)
@@
Effect of different knowledge reasoning methods of Answerer.
In Tab. 3, we present a comparison of Answerer using different knowledge reasoning methods. The results show that the performance using the strategy of voting surpasses that of concatenating under different knowledge selection settings. We argue that directly combining all the knowledge documents into a lengthened document makes it difficult for Answerer to reason on them, which is easily influenced by noisy information. In contrast, it is easier for Answerer to reason on each document to predict the answer. Simple voting can choose the best answer.
Table 3: Effect of different knowledge reasoning methods of Answerer. Concatenating denotes that we combine the key knowledge documents into one sentence and feed it into Answerer to predict the final answer. Voting means that we feed different key knowledge documents into Answerer to predict different answers and choose the best answer based on majority voting.
Method VQA Model Acc (%)
Concatenating BLIP2 (fine-tuned) 59.11
@@
Concatenating Ours 62.06
@@
Effect of the self-bootstrap learning method.
To evaluate the effectiveness of our self-bootstrap learning method, we compare the method with the strategy of independent training of two modules. We finetune the Answerer with the knowledge documents retrieved by DPR as the baseline. Independent training means that we take two passes in answerer training and one pass for selector training on the entire dataset. Cycle training means that we train the answerer and selector on each batch data of the dataset simultaneously. The results in Tab. 4 show that the model with cycle training outperforms the model with independent training by 3.81%. The VQA score of using independent training is even lower than the baseline. These results demonstrate that our cycle training method can effectively boost the Selector and Answerer each other, which makes the model find key knowledge documents and leverage the knowledge to answer questions.
Table 4: Effect of the self-bootstrap learning method.
Method Acc (%)
@@
Effect of different methods of pseudo-labeling.
In Tab. 5, we compare the model performance with different methods of pseudo-labeling. When using the model predictions as guidance, the model has a VQA score of 62.31%. When adding the weak supervision as the guidance, the model’s VQA score increases from 62.31% to 62.83%. The results demonstrate that using weak supervision labels preserves potentially useful documents, aiding the Answerer in accurately answering questions.
Table 5: Ablation study on different methods of pseudo-labeling.
Model predictions Weak supervision labels Acc (%)
@@
Effect of key knowledge documents selection ranges and quantities.
In Tab. 6, we evaluate key knowledge document selection using various numbers of candidate documents and selected documents. From the results, we have the following findings: (1) As the number of selected documents increases, the model’s performance improves. This indicates that using more documents to train and test contributes to answering questions. (2) Using more documents for training can improve the performance a lot (the 2nd line $v.s.$ the last line). However, using more documents for testing has almost no improvement (the 3rd line $v.s.$ 4th line). (3) When the number of candidate documents increases, the model’s performance improves. The result demonstrates that low-ranked documents based on DPR scores may contain useful information for question answering. It is necessary for the model to select key knowledge documents.
Table 6: Ablation study on different numbers of candidate documents and selected documents.
$K_{candidate}$ $K_{train}$ $K_{test}$ Acc (%)
@@
Table 7: Ablation study on different documents selection in Answerer fine-tuning.
Knowledge Selection Acc (%)
@@
Figure 2: Qualitative results on the test split of OK-VQA. We compared our method with a model that fine-tunes BLIP2 with knowledge ranked by DPR. The middle segment of the graph represents knowledge from various methods used to answer questions. On the right side of the graph, different answers are depicted when using distinct knowledge. Green and red colors indicate whether the selected final answer is correct.
Effect of different knowledge documents selection in Answerer fine-tuning.
Tab. 7 compares Answerer fine-tuning with different document selection strategies. The results show that our framework performs optimally when utilizing Selector in both Answerer training and inference. This is likely because the Selector provides more informative key knowledge documents and using both Selector ensures the consistency between the training domain and testing domain.
Performance of the knowledge retrieval.
In tab. 8, we evaluate our Selector in the knowledge retrieval task. Following previous methods Luo et al. (2021); Lin and Byrne (2022), we adopt pseudo relevance to measure if the retrieved document is relevant to the query due to the absence of ground-truth document. We use Recall to measure the performance of the the knowledge retrieval. From the results, we can see that our Selector improves the performance of DPR a lot. This means our Selector can retrieve more relevant knowledge documents which help answer questions. Compared to other retrievers, our Selector achieves the second best performance. Although FLMR outperforms our Selector in the knowledge retrieval, our framework achieves better accuracy in VQA (shown in Tab. 1) with the same backbone. This indicates that the knowledge documents selected by Selector have better consistency with Answerer.
Table 8: Retrieval performance on Google Search (GS).
Retriever R@5 R@10
VRR Luo et al. (2021) 80.4 88.55
RA-VQA-FrDPR Lin and Byrne (2022) 81.25 88.51
RA-VQA Lin and Byrne (2022) 82.84 89.00
FLMR Lin et al. (2023) 89.32 94.02
DPR Lin and Byrne (2022) 82.93 89.95
Our Selector 88.66 93.56
4.4 Qualitative Analysis